OpenStudy (anonymous):

What are the x-coordinates for the maximum points in the function f(x)=4cos(2x-pi) from x=0 to x=2pi

OpenStudy (freckles):

Well isn't the range of y=cos(u) is [-1,1] So the max value for y=cos(u) is 1 And the min value for y=cos(u) is -1 -- So the max value for f(x)=4cos(u) is?

OpenStudy (freckles):

f(u)=4cos(u) *

OpenStudy (anonymous):

I'm not quite sure if I follow

OpenStudy (freckles):

cos(u) has max value 1 so 4cos(u) has max value 4(1)=4

OpenStudy (freckles):

Does that make sense?

OpenStudy (anonymous):

Yes

OpenStudy (freckles):

So the max value of 4cos(2x-pi) is 4(1)=4 so you need to solve 4cos(2x-pi)=4 to find for which x 4cos(2x-pi) is 4

OpenStudy (freckles):

Or basically solve cos(2x-pi)=1

OpenStudy (ribhu):

x = pi/3

OpenStudy (ribhu):

sorry x=pi/2

OpenStudy (freckles):

should be two values

OpenStudy (anonymous):

And x=3pi/2?

OpenStudy (freckles):

yep

OpenStudy (ribhu):

correct @Patsfan12 @freckles

OpenStudy (anonymous):

Alright thanks guys!